OpenStudy (thomas5267):

This is an assignment so don't give me the answer, only hints. Could someone interpret this question for me? Write the elements of \(\mathbb{R}^n\) and \(\mathbb{R}^m\) as rows. If \(A\) is an \(m\times n\) matrix, define \(T\colon\mathbb{R}^m\rightarrow \mathbb{R}^n\) by \(T(\mathbf{y})=\mathbf{y}A\) for all rows \(\mathbf{y}\) in \(\mathbb{R}^n\). Show that the rows of \(A\) are \(T(\mathbf{f}_1),\ldots,T(\mathbf{f}_m)\) where \(\mathbf{f}_i\) denotes row \(i\) of \(I_m\).

OpenStudy (thomas5267):

The domain of \(T\) is \(\mathbb{R}^m\) but \(\mathbf{y}\in\mathbb{R}^n\)?

OpenStudy (ikram002p):

since http://prntscr.com/455j8w

OpenStudy (thomas5267):

\(\mathbf{y}\in\mathbb{R}^{\color{red} n}\)

OpenStudy (ikram002p):

as u can see T(y) = y A defined on R^n to R^n then y in R^n y A in R^n

OpenStudy (thomas5267):

\(T\colon\mathbb{R}^{\color{red} m}\rightarrow\mathbb{R}^n\)

OpenStudy (ikram002p):

well its given :O if u have T : C^n to R^n then y in C^n

OpenStudy (ikram002p):

wait i got u nw the qn R^m to R^n nw i got what u mean so why y in R^m also right ?

OpenStudy (ikram002p):

well consider this R^m is a group like ? so its closed , means if ab=c s.t c in R^m then a in R^m and b in R^m

OpenStudy (ikram002p):

its not wrong :D

OpenStudy (thomas5267):

Is there a typo because \(A_{m\times n}\), \(T\colon\:\mathbb{R}^{\color{red}m}\rightarrow\mathbb{R}^n,\,T(\mathbf{y})=\mathbf{y}A\) yet \(\mathbf{y}\in\mathbb{R}^{\color{red}n}\)?

OpenStudy (ikram002p):

i dnt think so , maybe its m for a reason

OpenStudy (anonymous):

@ikram002p I think we should start at f(i) is row of I_m. f(i) is in R^m T(I_m)=I_m*A

OpenStudy (ikram002p):

so no typo ?

OpenStudy (thomas5267):

But how could you multiply a 1xn matrix with a mxn matrix?

OpenStudy (ikram002p):

u cant x_X

OpenStudy (ikram002p):

got what u mean , a bit confused sry :| for definition of grouping no problem but solving with matrix ,,, mm